I would like the answer to part B as I have already completed A. Please also indicate why the Price Discount of $10 is not a Variable Cost and does not affect the Contribution Margin i.e decrease the contribution margin from 20 to 10.

Company X manufactures soft drinks. Its manufacturing plant has the capacity to produce 10,000 cases each month; current production and sales are 7,500 cases per month. The company normally charges $150 per case. Cost information for the current activity level is as follows:

Variable costs that vary with units produced

Direct materials ---------------------------------------------$ 262,500

Direct manufacturing labour -------------------------------- 300,000

Variable costs (for setups, materials handling, quality control, and so on) that vary with number of batches, 150 batches x $500 per batch ---------------------------------------75,000 Fixed manufacturing costs ------------------------------------275,000

Fixed marketing costs -----------------------------------------175,000

Total costs ---------------------------------------------------$ 1,087,500

Company X has just received a special one-time-only order for 2,500 cases at $100 per case. Company X usually makes its soft drinks for its existing customers in batch sizes of 50 case (150 batches x 50 cases per batch = 7,500 cases). The special order requires them to make the cases in 25 batches of 100 each

(a) Should Company X accept this special order? Why? Explain briefly.

PARTICULARS

CURRENT CALCULATION

CURRENT

SPECIAL ORDER CALCULATION

SPECIAL ORDER

Direct material per case

262,500/7,500

35

35

Direct Labour per case

300,000/,7500

40

40

Variable cost

$500 × 150

     7,500

10

$500 × 25

     2,500

5

Total Variable Cost (Relevant)

35 + 40 + 10 = 85

35 + 40 + 5 = 80

Selling Price

7500 × $150 = $1,125,000

150

2500 × $100 = $25,000

100

Contribution Margin

150 – 85 = 65

100 – 80 = 20

Total additional Contribution

20 × 2500 cases

Net Income

$50,000

$50,000

The offer should be accepted.

(b) As in requirement (a) assume that monthly capacity is 10,000 cases. Company X is concerned that if it accepts the special order, its existing customers will immediately demand a price discount of $10 in the month in which the special order is being filled. They would argue that Company X capacity costs are now being spread over more units, and that existing customers should get the benefit of these lower costs. Should Company X accept the special order under these conditions? Show all calculations.
Answer #1
(b) If the order is accepted , Sales price per unit for 75000 cases will reduce by $10
There will be total   decrease in revenue by (7500*10)=$75000fromexisting customers
PARTICULARS CURRENT CALCULATION CURRENT SPECIAL ORDER CALCULATION SPECIAL ORDER
Direct material per case 262,500/7,500 35 35
Direct Labour per case 300,000/,7500 40 40
Variable cost $500 × 150 10 $500 × 25 5
     7,500      2,500
Total Variable Cost (Relevant) 35 + 40 + 10 = 85 35 + 40 + 5 = 80
Selling Price 7500 × $150 = $1,125,000 150 2500 × $100 = $25,000 100
Contribution Margin 150 – 85 = 65 100 – 80 = 20
Total additional Contribution 20 × 2500 cases
Net Income $50,000 $50,000
Total Additional contribution without reduction in price for existing customers $50,000
Reduction in contribution if reduction in price for existing customers is allowed $75,000 (7500*10)
Net additional contribution from the new order ($25,000) (50000-75000)
Since the new order will reduce profit by $25000, This order should not be accepted
Note:
Reduction in price will reduce the contribution per unit from existing customers
The alternative method of solving this problem is to find the current contribution from existing customers
Then calculate the contribution from existing customers with reduction in price and additional contribution from the new order
Compare the two
The method shown above is simpler method